# RectiCast: Rectifying Distribution Shift in Cascaded Precipitation Nowcasting

### v2

@@ Abstract
Precipitation nowcasting, which aims to provide high spatio-temporal resolution precipitation forecasts by leveraging current radar observations, is a core task in regional weather forecasting. Recently, the cascaded architecture has emerged as the mainstream paradigm for deep learning-based precipitation nowcasting. This paradigm involves a deterministic model to predict posterior mean, followed by a probabilistic model to generate local stochasticity. However, existing methods commonly overlook the conflation of the systematic distribution shift in deterministic predictions and the local stochasticity. As a result, the distribution shift of the deterministic component contaminates the predictions of the probabilistic component, leading to inaccuracies in precipitation patterns and intensity, particularly over longer lead times. To address this issue, we introduce RectiCast, a two-stage framework that explicitly decouples the rectification of mean-field shift from the generation of local stochasticity via a dual Flow Matching model. In the first stage, a deterministic model generates the posterior mean. In the second stage, we introduce a Rectifier to explicitly learn the distribution shift and produce a rectified mean. Subsequently, a Generator focuses on modeling the local stochasticity conditioned on the rectified mean. Experiments on two radar datasets demonstrate that RectiCast achieves significant performance improvements over existing state-of-the-art methods.
@@ 1 Introduction
Precipitation nowcasting, which aims to provide high spatio-temporal resolution forecasts for up to 2 hours based on current observations, is a core task in regional weather forecasting. Conventional deep learning methods are primarily categorized into deterministic and probabilistic models. Deterministic models directly optimize the mean squared error (MSE) against the ground-truth, effectively capturing macroscopic motion trends. However, their predictions lack fine-grained local details. Probabilistic models model the entire precipitation system stochastically. While capable of generating predictions with realistic local details, the predicted location of precipitation is often inaccurate [diffcast]. Recently, to integrate the strengths of both approaches, several works [diffcast, percpcast, cascast] have employed a cascaded architecture. This framework treats the deterministic prediction as the posterior mean (i.e., the macroscopic trend) given the observation sequence [diffcast, percpcast]. A probabilistic model then generates the residual between the ground-truth distribution and this mean (i.e., the local stochasticity) [diffcast]. This paradigm has proven effective in capturing the evolution of weather systems at multiple scales.
@@ 2 Related Work
Deep learning-based precipitation nowcasting methods are generally categorized into deterministic, probabilistic, and cascaded approaches. Deterministic models excel at capturing the overall trend of precipitation movement. ConvLSTM [convlstm] combines convolutional operators with recurrent units to extract spatio-temporal features, leading to variants such as TrajGRU [trajgru] and PredRNN [predrnn]. PhyDNet [phydnet] decomposes the prediction into PDE-guided physical motion and random motion to enhance physical consistency. SimVP [simvp] employs a simplified convolutional encoder-decoder architecture for greater computational efficiency. Earthformer [earthformer] and Earthfarseer [earthfarseer] introduce Transformer [transformer] to precipitation nowcasting. However, deterministic models tend to predict an average of multiple future scenarios, leading to forecasts that become blurry over time. Probabilistic models generate more realistic details by modeling the data distribution of future precipitation, but they often fail to accurately capturing large-scale precipitation motion. DGMR [dgmr] introduces spatial and temporal discriminators to constrain the prediction distribution through adversarial training. NowcastNet [nowcastnet] introduces physical constraints to regularize the forecast of generative backbone. PreDiff [prediff] constructs a diffusion model in the latent space and designs a guidance module to adjust the denoising process. MCVD [mcvd] learns the inter-frame relationships of precipitation through random masking. Cascaded models combine the strengths of the previous two approaches, significantly improving forecast accuracy. DiffCast [diffcast] introduces a residual diffusion mechanism to model local stochastic variations conditioned on a deterministic forecast. CasCast [cascast] performs the diffusion process in latent space, generating the entire forecast sequence at once. PercpCast [percpcast] employs a perceptual loss to constrain the generation process, achieving strong results in long-range forecasting.
@@ 3.2 Preliminary: Flow Matching
Flow Matching [flow] is an emerging generative framework. Its core idea is to learn the conditional velocity along a linear path from pure Gaussian noise to a real data sample. Let $f_{\theta}$ denote the generative model, $\epsilon\sim\mathcal{N}(0,I)$ be Gaussian noise, $x$ be a real data sample from the training set, and $t\sim U(0,1)$ be a timestep uniformly sampled from $[0,1]$. The intermediate state $x_{t}$ on this linear path is defined as:
@@ 4.1 Deterministic Predictive Backbone
In the first stage of prediction, we use a deterministic predictive backbone $D_{\theta_{1}}$ to predict the posterior mean $\mu^{\text{raw}}$. Similar to [diffcast], our framework supports any prediction models of recurrent-based and recurrent-free architectures. We train the deterministic predictive backbone to minimize the mean-squared error (MSE) between the posterior mean and the ground truth $y$:
@@
This enables $D_{\theta_{1}}$ to model the distribution of $p_{\theta_{1}}(\mu^{\text{raw}}|x)$, which is the distribution of the posterior mean of $y$ given the current observation $x$. The prediction $\mu^{\text{raw}}$ obtained this way effectively captures the global motion trend and serves as a strong denoising prior for the subsequent probabilistic component. However, it still suffers from distribution shifts over time.
@@ 4.2 Rectifier-Generator Framework
As mentioned, conditioning the probabilistic model directly on the raw prediction $\mu^{\text{raw}}$ conflates its learning objectives. This approach forces the model to disentangle the distribution shift of $\mu^{\text{raw}}$ and the local stochastic process of precipitation from a single entangled distribution. To explicitly disentangle the distribution shift from the local stochasticity, we train two separate Flow Matching models in the second stage to model each component.
The first model, the Rectifier ($\mathcal{R}_{\theta_{2}}$), models the distribution shift of the raw posterior mean sequence $\mu^{\text{raw}}$ over lead time. While a posterior mean entirely free of distribution shift is unattainable, the shift in predictions with shorter lead times remains within an acceptable range for modeling local stochasticity. Consequently, we utilize the initial segment of $\mu^{\text{raw}}$ as the target distribution for $\mathcal{R}_{\theta_{2}}$. Following the optimal configuration in [diffcast], the segment length $m$ is set equal to the input sequence length. We model the rectification process autoregressively via a Spatial-Temporal U-Net equipped with a side encoder, which will be discussed in 4.3. We define the target for the $i$-th segment as $\mu^{\text{rec}}_{i}=D_{\theta_{1}}(s_{i-1})_{1}$, for $i=2,3,\dots,\lceil L_{\text{out}}/m\rceil$, where $D_{\theta_{1}}(s_{i-1})_{1}$ represents the first segment of the backbone prediction given the previous ground truth segment $s_{i-1}$. For the first segment ($i=1$), no rectification applies, i.e., $\mu^{\text{rec}}_{1}=\mu^{\text{raw}}_{1}$. The Rectifier is conditioned on the current raw mean segment $\mu^{\text{raw}}_{i}$ and the previous rectified mean segment $\mu^{\text{rec}}_{i-1}$. In addition to the timestep $t$, the network takes the segment index $i$ as a condition to indicate the degree of shift. For the $i$-th segment, $\mathcal{R}_{\theta_{2}}$ establishes the following conditional distribution:
@@
The second flow model, the Generator ($\mathcal{G}_{\theta_{3}}$), focuses on modeling local stochasticity. It employs the same network architecture as $\mathcal{R}_{\theta_{2}}$ and forecasts autoregressively. $\mathcal{G}_{\theta_{3}}$ is conditioned on the prediction from the previous segment $\hat{y}_{i-1}$ and the current rectified mean $\mu^{\text{rec}}_{i}$. For the $i$-th segment, $\mathcal{G}_{\theta_{3}}$ models:
@@ 4.3 Network Architecture
This subsection details the architecture of the Rectifier and Generator. The architectures of both the Rectifier and Generator consist of a U-Net backbone and a side encoder. The structure of the side encoder is similar to that of the U-Net encoder. Each Spatio-Temporal Block in the U-Net and the side encoder is composed of several spatial residual blocks and a temporal attention block. The spatial residual blocks independently extract spatial features from each input frame. The temporal attention block employs bottleneck point-wise convolutions and channel attention mechanisms to model inter-frame dependencies. Conditions for Flow Matching are injected through both the backbone and the side encoder. The backbone condition is concatenated with $z_{t}$ frame-by-frame along the channel dimension. The side-injected condition is fed into the side encoder to generate scale and shift parameters, which are then injected via FiLM [film] before the spatial residual blocks at the corresponding scales in the U-Net. To prevent the model from learning shortcuts, the conditioning from the side network is injected only into deeper layers of the U-Net backbone.
@@ 5.1 Experimental Setting
Dataset: Experiments are conducted on SEVIR [sevir] and MeteoNet [meteonet]. SEVIR contains 20,394 sequences of radar frames for weather events from 2017 to 2020. Each radar sequence covers a 384 km $\times$ 384 km area with a 1 km resolution, comprising 49 frames captured every 5 minutes over a 4-hour period. We split the dataset into training, validation, and test sets with the time points January 1, 2019, and July 1, 2019, respectively. Samples of 25 consecutive frames are extracted from each sequence with a stride of 5 frames. MeteoNet contains precipitation radar data from northwestern and southeastern France from 2016 to 2018. The radar sequences cover a 550 km $\times$ 550 km area with a 5-minute observation interval. Dataset splitting and anomaly filtering of MeteoNet follow [diffcast]. The spatial dimensions of all datasets are downsampled to $\times128$. The prediction task is to forecast the next 20 frames given 5 initial frames.
Evaluation: We evaluate the model’s nowcasting performance using the average Critical Success Index (CSI) and Heidke Skill Score (HSS), calculated as in [prediff, recon, sevir]. CSI measures pixel-wise matching between predictions and ground truth over a set of thresholds, while HSS measures the forecast skill improvement relative to random forecast. We follow [diffcast] for data scaling and threshold selection on both SEVIR and MeteoNet datasets. To evaluate forecast accuracy at different spatial scales, we also compute CSI on predictions downsampled via $\times4$ and $\times16$ max-pooling following [diffcast]. Additionally, we calculate the Structural Similarity (SSIM) and the Learned Perceptual Image Patch Similarity (LPIPS) to evaluate the visual quality of the forecasts.
Baselines: Seven notable models are selected as baselines, including three deterministic models, one probabilistic model, and three cascaded models. ConvLSTM [convlstm], SimVP [simvp], and Earthformer [earthformer] are deterministic models based on RNN, CNN, and Transformer, respectively. Prediff [prediff] is a probabilistic model based on the Latent Diffusion Model. DiffCast [diffcast], CasCast [cascast], and PercpCast [percpcast] are recent state-of-the-art cascaded models.
Implementation Details: All models were trained using the Adam optimizer with a cosine learning rate scheduler. The maximum and minimum learning rates were $10^{-4}$ and $10^{-7}$, respectively. We trained the models for 200k iterations on SEVIR and 80k iterations on MeteoNet. All cascaded models were trained in two stages. To ensure a fair comparison of the probabilistic components, SimVP [simvp] was used uniformly as the deterministic component in all cascaded models. During inference, diffusion-based and flow-based modules employ the DDIM sampler [ddim] and the Euler solver, respectively. The number of sampling steps following the experimental settings in their respective original papers. All experiments were conducted on 4 V100 GPUs with a global batch size of 24.
@@ 5.2 Experimental Results
To investigate how the decoupled learning architecture improves the forecasting capability of RectiCast, we plot the three CSI curves of all cascaded models and their deterministic component SimVP [simvp] against lead time in Figure 3. As the lead time increases, prediction uncertainty grows, causing the performance of all models to decline. However, the performance of different cascaded models declines at different rates. This suggests that the cascaded structures are affected differently by the distribution shift of the deterministic prediction. PercpCast [percpcast] performs best at short lead times, but its performance drops sharply as SimVP’s performance declines. This is because its probabilistic part directly models a frame-wise mapping from the deterministic prediction to the final forecast, making it the most susceptible to the distribution shift of the deterministic prediction. While DiffCast [diffcast] and CasCast [cascast] do not explicitly model the distribution shift of the deterministic prediction, their modeling of inter-frame dependencies allows them to maintain better performance at longer lead times. RectiCast’s three CSI curves decline the most slowly over time, suggesting that its explicit modeling of the deterministic prediction’s distribution shift significantly enhances its long-term forecasting capability. This again validates the effectiveness of RectiCast’s Rectifier-Generator framework.
Figure 4: A case study comparing forecasts from different cascaded models on the SEVIR dataset.
@@ 5.3 Ablation Study
To verify the validity of the individual components in RectiCast, we conducted ablation experiments on the SEVIR dataset. We compared the nowcasting performance across several configurations: the full model with both the Rectifier and Generator, a variant without the Generator, and variants without the Rectifier. Without the Generator, the model degenerates into a pure mean rectifier. When the Rectifier is removed, we report metrics for two scenarios: the Generator learning the precipitation distribution $y$ directly, and learning the residual between the distribution and the raw mean $y-\mu^{\text{raw}}$. As shown in Table 2, ablating either the Rectifier or the Generator leads to a significant performance drop. This indicates that without the Generator, the model fails to generate precise local details. Moreover, the absence of the Rectifier causes the mean shift to interfere with the learning of local stochastics. These results confirm the effectiveness of decoupling the learning of the mean shift and local stochastics.
Finally, we analyze the computational efficiency in Table 3. Although RectiCast entails a higher Number of Function Evaluations (NFE) due to its dual Flow Matching architecture and autoregressive manner, it maintains a fast inference speed. This validates that our framework achieves a favorable trade-off between forecast quality and computational cost.

### v3

@@ Abstract
Precipitation nowcasting, which aims to provide high spatio-temporal resolution precipitation forecasts by leveraging current radar observations, is a core task in regional weather forecasting. Recently, the cascaded architecture has emerged as the mainstream paradigm for deep learning-based precipitation nowcasting. This paradigm involves a deterministic model to predict posterior mean, followed by a probabilistic model to generate local stochasticity. However, existing methods commonly conflate the systematic distribution shift in deterministic predictions with local stochasticity, which contaminates the probabilistic component and degrades forecast accuracy, especially over longer lead times. To address this issue, we introduce RectiCast, a two-stage framework that explicitly decouples the rectification of mean-field shift from the generation of local stochasticity via a dual Flow Matching model. In the first stage, a deterministic model generates the posterior mean. In the second stage, a Rectifier explicitly learns the distribution shift to yield a rectified mean, conditioning a Generator that models local stochasticity. Experiments on two radar datasets demonstrate that RectiCast achieves significant performance improvements over existing state-of-the-art methods.
@@ 1 Introduction
Precipitation nowcasting, which aims to provide high spatio-temporal resolution forecasts for up to 2 hours based on current observations, is a core task in regional weather forecasting. Conventional deep learning methods are primarily categorized into deterministic and probabilistic models. Deterministic models directly optimize the mean squared error (MSE) against the ground-truth, effectively capturing macroscopic motion trends. However, their predictions lack fine-grained local details. Probabilistic models model the entire precipitation system stochastically. While capable of generating predictions with realistic local details, the predicted location of precipitation is often inaccurate [20]. Recently, to integrate the strengths of both approaches, several works [20, 1, 5] have employed a cascaded architecture. This framework treats the deterministic prediction as the posterior mean (i.e., the macroscopic trend) given the observation sequence [20, 1]. A probabilistic model then generates the residual between the ground-truth distribution and this mean (i.e., the local stochasticity) [20]. This paradigm has proven effective in capturing the evolution of weather systems at multiple scales.
@@ 2 Related Work
Deep learning-based precipitation nowcasting methods are generally categorized into deterministic, probabilistic, and cascaded approaches. Deterministic models excel at capturing the overall trend of precipitation movement. ConvLSTM [12] combines convolutional operators with recurrent units to extract spatio-temporal features, leading to variants such as TrajGRU [13] and PredRNN [18]. PhyDNet [7] decomposes the prediction into PDE-guided physical motion and random motion to enhance physical consistency. SimVP [3] employs a simplified convolutional encoder-decoder architecture for greater computational efficiency. Earthformer [4] and Earthfarseer [19] introduce Transformer [15] to precipitation nowcasting. However, deterministic models tend to predict an average of multiple future scenarios, leading to forecasts that become blurry over time. Probabilistic models generate more realistic details by modeling the data distribution of future precipitation, but they often fail to accurately capture large-scale precipitation motion. DGMR [11] introduces spatial and temporal discriminators to constrain the prediction distribution through adversarial training. NowcastNet [21] introduces physical constraints to regularize the forecast of generative backbone. PreDiff [2] constructs a diffusion model in the latent space and designs a guidance module to adjust the denoising process. MCVD [17] learns the inter-frame relationships of precipitation through random masking. Cascaded models combine the strengths of the previous two approaches, significantly improving forecast accuracy. DiffCast [20] introduces a residual diffusion mechanism to model local stochastic variations conditioned on a deterministic forecast. CasCast [5] performs the diffusion process in latent space, generating the entire forecast sequence at once. PercpCast [1] employs a perceptual loss to constrain the generation process, achieving strong results in long-range forecasting.
@@ 3.2 Preliminary: Flow Matching
Flow Matching [8] is an emerging generative framework. Its core idea is to learn the conditional velocity along a linear path from pure Gaussian noise to a real data sample. Let $f_{\theta}$ denote the generative model, $\epsilon\sim\mathcal{N}(0,I)$ be Gaussian noise, $x$ be a real data sample from the training set, and $t\sim U(0,1)$ be a timestep uniformly sampled from $[0,1]$. The intermediate state $x_{t}$ on this linear path is defined as:
@@ 4.1 Deterministic Predictive Backbone
In the first stage of prediction, we use a deterministic predictive backbone $D_{\theta_{1}}$ to predict the posterior mean $\mu^{\text{raw}}$. Similar to [20], our framework supports any prediction models of recurrent-based and recurrent-free architectures. We train the deterministic predictive backbone to minimize the mean-squared error (MSE) between the posterior mean and the ground truth $y$:
@@
This enables $D_{\theta_{1}}$ to model the distribution of $p_{\theta_{1}}(\mu^{\text{raw}}|x)$, which is the distribution of the posterior mean of $y$ given the current observation $x$. The prediction $\mu^{\text{raw}}$ obtained this way effectively captures the global motion trend and serves as a strong denoising prior for the subsequent probabilistic component.
@@ 4.2 Rectifier-Generator Framework
As mentioned, conditioning the probabilistic model directly on $\mu^{\text{raw}}$ conflates its learning objectives. To explicitly disentangle the distribution shift from the local stochasticity, we train two separate Flow Matching models in the second stage to model each component.
The first model, the Rectifier $\mathcal{R}_{\theta_{2}}$, models the distribution shift of the raw posterior mean sequence $\mu^{\text{raw}}$ over lead time. While a posterior mean entirely free of distribution shift is unattainable, the shift in predictions with shorter lead times remains within an acceptable range for modeling local stochasticity. Consequently, we utilize the initial segment of $\mu^{\text{raw}}$ as the target distribution for $\mathcal{R}_{\theta_{2}}$. Following the optimal configuration in [20], the segment length $m$ is set equal to the input sequence length. We model the rectification process autoregressively via a Spatial-Temporal U-Net equipped with a side encoder, which will be discussed in 4.3. We define the target for the $i$-th segment as $\mu^{\text{rec}}_{i}=D_{\theta_{1}}(s_{i-1})_{1}$, for $i=2,3,\dots,\lceil L_{\text{out}}/m\rceil$, where $D_{\theta_{1}}(s_{i-1})_{1}$ represents the first segment of the backbone prediction given the previous ground truth segment $s_{i-1}$. For the first segment ($i=1$), no rectification applies, i.e., $\mu^{\text{rec}}_{1}=\mu^{\text{raw}}_{1}$. The Rectifier is conditioned on the current raw mean segment $\mu^{\text{raw}}_{i}$ and the previous rectified mean segment $\mu^{\text{rec}}_{i-1}$. In addition to the timestep $t$, the network takes the segment index $i$ as a condition to indicate the degree of shift. For the $i$-th segment, $\mathcal{R}_{\theta_{2}}$ establishes the following conditional distribution:
@@
The second model, the Generator $\mathcal{G}_{\theta_{3}}$, focuses on modeling local stochasticity. It employs the same network architecture as $\mathcal{R}_{\theta_{2}}$ and forecasts autoregressively. $\mathcal{G}_{\theta_{3}}$ is conditioned on the prediction from the previous segment $\hat{y}_{i-1}$ and the current rectified mean $\mu^{\text{rec}}_{i}$. For the $i$-th segment, $\mathcal{G}_{\theta_{3}}$ models:
@@ 4.3 Network Architecture
This subsection details the architecture of the Rectifier and Generator. The architectures of both the Rectifier and Generator consist of a U-Net backbone and a side encoder. The structure of the side encoder is similar to that of the U-Net encoder. Each Spatio-Temporal Block in the U-Net and the side encoder is composed of several spatial residual blocks and a temporal attention block. The spatial residual blocks independently extract spatial features from each input frame. The temporal attention block employs bottleneck point-wise convolutions and channel attention mechanisms to model inter-frame dependencies. Conditions for Flow Matching are injected through both the backbone and the side encoder. The backbone condition is concatenated with $z_{t}$ frame-by-frame along the channel dimension. The side-injected condition is fed into the side encoder to generate scale and shift parameters, which are then injected via FiLM [10] before the spatial residual blocks at the corresponding scales in the U-Net. To prevent the model from learning shortcuts, the conditioning from the side network is injected only into deeper layers of the U-Net backbone.
@@ 5.1 Experimental Setting
Dataset: Experiments are conducted on SEVIR [16] and MeteoNet [6]. SEVIR contains 20,394 sequences of radar frames for weather events from 2017 to 2020. Each radar sequence covers a 384 km $\times$ 384 km area with a 1 km resolution, comprising 49 frames captured every 5 minutes over a 4-hour period. We split the dataset into training, validation, and test sets with the time points January 1, 2019, and July 1, 2019, respectively. Samples of 25 consecutive frames are extracted from each sequence with a stride of 5 frames. MeteoNet contains precipitation radar data from northwestern and southeastern France from 2016 to 2018. The radar sequences cover a 550 km $\times$ 550 km area with a 5-minute observation interval. Dataset splitting and anomaly filtering of MeteoNet follow [20]. The spatial dimensions of all datasets are downsampled to $\times128$. The prediction task is to forecast the next 20 frames given 5 initial frames.
Evaluation: We evaluate the model’s nowcasting performance using the average Critical Success Index (CSI) and Heidke Skill Score (HSS), calculated as in [2, 9, 16]. CSI measures pixel-wise matching between predictions and ground truth over a set of thresholds, while HSS measures the forecast skill improvement relative to random forecast. We follow [20] for data scaling and threshold selection on both SEVIR and MeteoNet datasets. To evaluate forecast accuracy at different spatial scales, we also compute CSI on predictions downsampled via $\times4$ and $\times16$ max-pooling following [20]. Additionally, we calculate the Structural Similarity (SSIM) and the Learned Perceptual Image Patch Similarity (LPIPS) to evaluate the visual quality of the forecasts.
Baselines: Seven notable models are selected as baselines, including three deterministic models, one probabilistic model, and three cascaded models. ConvLSTM [12], SimVP [3], and Earthformer [4] are deterministic models based on RNN, CNN, and Transformer, respectively. Prediff [2] is a probabilistic model based on the Latent Diffusion Model. DiffCast [20], CasCast [5], and PercpCast [1] are recent state-of-the-art cascaded models.
Implementation Details: All models were trained using the Adam optimizer with a cosine learning rate scheduler. The maximum and minimum learning rates were $10^{-4}$ and $10^{-7}$, respectively. We trained the models for 200k iterations on SEVIR and 80k iterations on MeteoNet. All cascaded models were trained in two stages. To ensure a fair comparison of the probabilistic components, SimVP [3] was used uniformly as the deterministic component in all cascaded models. During inference, diffusion-based and flow-based modules employ the DDIM sampler [14] and the Euler solver, respectively. The number of sampling steps following the experimental settings in their papers. All experiments were conducted on 4 V100 GPUs with a global batch size of 24.
@@ 5.2 Experimental Results
To investigate how the decoupled learning architecture improves the forecasting capability of RectiCast, we plot the three CSI curves of all cascaded models and their deterministic component SimVP [3] against lead time in Figure 3. As the lead time increases, prediction uncertainty grows, causing the performance of all models to decline. However, the performance of different cascaded models declines at different rates. This suggests that the cascaded structures are affected differently by the distribution shift of the deterministic prediction. PercpCast [1] performs best at short lead times, but its performance drops sharply as SimVP’s performance declines. This is because its probabilistic part directly models a frame-wise mapping from the deterministic prediction to the final forecast, making it the most susceptible to the distribution shift of the deterministic prediction. While DiffCast [20] and CasCast [5] do not explicitly model the distribution shift of the deterministic prediction, their modeling of inter-frame dependencies allows them to maintain better performance at longer lead times. RectiCast’s three CSI curves decline the most slowly over time, suggesting that its explicit modeling of the deterministic prediction’s distribution shift significantly enhances its long-term forecasting capability. This again validates the effectiveness of RectiCast’s Rectifier-Generator framework.
Figure 4: A case study from different cascaded models on the SEVIR dataset.
@@ 5.3 Ablation Study
To verify the validity of the individual components in RectiCast, we conducted ablation experiments on the SEVIR dataset. We compared the nowcasting performance across several configurations: the full model with both the Rectifier and Generator, a variant without the Generator, and variants without the Rectifier. Without the Generator, the model degenerates into a pure mean rectifier. When the Rectifier is removed, we report metrics for two scenarios: the Generator learning the precipitation distribution $y$ directly, and learning the residual between the distribution and the raw mean $y-\mu^{\text{raw}}$. As shown in Table 2, ablating either the Rectifier or the Generator leads to a significant performance drop. This indicates that without the Generator, the model fails to generate precise local details. Moreover, the absence of the Rectifier causes the mean shift to interfere with the learning of local stochastics. These results confirm the effectiveness of decoupling the learning of the distribution shift and local stochastics.
Finally, we analyze the computational efficiency in Table 3. Although RectiCast’s autoregressive manner and dual Flow Matching architecture increases the Number of Function Evaluations (NFE), it maintains fast inference speed, validating a favorable accuracy-efficiency trade-off.